# Compute Optimal Inference and Provable Amortisation Gap in Sparse Autoencoders

### v1

@@ 4 Synthetic Sparse Inference Experiments
We present the results of our experiments comparing different sparse encoding strategies across various scenarios. To provide a minimal setting for investigating the phenomena of interest, all experiments were conducted using synthetic data with $N=16$ sparse sources, $M=8$ measurements, and $K=3$ active components per timestep, unless otherwise specified (more settings in Sec. 4.4).
@@ 4.3 Unknown Sparse Codes and Dictionary
Finally, we evaluate all four methods when both latent representations and dictionary are unknown. We use a dataset of $2048$ samples, evenly split between training and testing sets, and conduct 5 independent runs of $100,000$ training steps each.
@@ 4.4 Performance Across Varying Data Regimes
We also investigated the trade-off between sparsity and performance for each method in Figure 6. Sparse coding achieves slightly lower reconstruction error for each L0 level, barring some very small active latents. Sparse coding shows a Pareto improvement at each L0 level in terms of MCC, even with very small active latents. The improvement is more evident when plotting against L1 rather than L0, as L1 accounts for the magnitude of non-zero values. The presence of very small non-zero latents in sparse coding motivates the exploration of top-$k$ sparse coding, detailed in Appendix F.2.
@@ 5 Interpretability of Sparse Coding Schemes
A common criticism of more powerful encoding techniques is that they can find concepts that are more difficult to interpret, or that are not actually used by the model. To investigate the interpretability of more complex encoding techniques, we trained both an SAE and MLP on $406$ million tokens from OpenWebText. The MLP used a hidden width of $4224$ and both models had $16,896$ sparse codes. The models were trained on the residual-stream pre-activations of Layer 9 in GPT-2 Small, which have a dimension of 768. We used a learning rate of $3\cdot 10^{-4}$ and an $L_{1}$ penalty of $1\cdot 10^{-4}$. Throughout training, we tracked normalised mean squared error (MSE), dividing it by the error when predicting the mean activations, as well as $L_{0}$ sparsity. The final SAE achieved a normalised MSE of $0.062$ with an average $L_{0}$ of $39.55$, while the MLP reached a normalised MSE of $0.051$ with an average $L_{0}$ of $40.20$. The final models had a significant number of dead neurons (measured as not having fired in the last $50,000$ training steps): $71\%$ for the SAE and $66\%$ for the MLP.
To assess the interpretability of the learned features, we selected $200$ random features that activated on our test set of $13.1$ million OpenWebText tokens. For each feature, we employed an automated interpretability classification approach. We identified the top $10$ activating examples for the feature in our test dataset and labelled the token with the highest activation. We also computed logit effects for each feature through the path expansion $W_{U}\cdot f$, where $W_{U}$ is the model’s unembedding matrix and $f$ is the feature vector. The top $10$ and bottom $10$ tokens resulting from this logit effect calculation were noted.
We provided both the activating examples and the top and bottom promoted logits to GPT-4, which was instructed to construct a precise explanation of the feature’s function (prompt in Appendix H). To evaluate the accuracy of these interpretations, we presented them to another instance of GPT-4 along with 5 new activating examples and 5 non-activating examples, labeling the token on which the feature potentially activates. The model was then asked to predict which of these examples the feature would activate on, and we calculated the F1-score compared to the ground truth.
Figure 7 displays the distributions of the F1 scores for the 200 SAE and MLP characteristics. The results indicate that the MLP features demonstrate interpretability comparable to the SAE features.
@@
Figure 7: Distribution of F1 scores for feature interpretability of an SAE and an MLP trained on residual stream activations of Layer 9 in GPT-2.
@@ 6 Discussion
Our study presents both theoretical and empirical evidence for the existence of an inherent amortisation gap in sparse autoencoders (SAEs) when applied to neural network interpretability tasks. We prove that SAEs with linear-nonlinear encoders cannot achieve optimal sparse recovery in settings where such recovery is theoretically possible. This limitation is corroborated by our experimental results, which demonstrate superior performance of more complex encoding methods, such as multilayer perceptrons (MLPs) and sparse coding, across various synthetic data scenarios. Notably, our investigation of GPT-2 activations reveals that MLP-based features exhibit interpretability comparable to, and in some cases exceeding, that of SAE features. These findings challenge the prevailing assumption that simpler encoders are necessary for maintaining interpretability. The results of our study have significant implications for the field of neural network interpretability, particularly in the context of large language models. They suggest that more sophisticated encoding techniques can potentially improve feature extraction and interpretability without compromising the validity of the extracted features. However, this potential improvement comes with increased computational costs. Our work provides a framework for exploring this trade-off.
The use of simple linear-nonlinear encoders in SAEs for language model interpretability has been primarily motivated by concerns that more powerful methods might extract features not actually utilised by the transformer (Bricken et al., 2023). However, this approach may be overly conservative given the complexity of representations in transformer layers, which result from multiple rounds of attention and feed-forward computations. Better encoders aligns with recent work on inference-time optimisation (Nanda et al., 2024), and will be validated as we improve encoding evaluation (Makelov et al., 2024). Regardless, SAEs are sensitive to hyperparameters and fragile (Cunningham et al., 2023), so exploring more powerful encoders is warranted.
The computational cost of more complex encoders should be weighed against potential benefits in feature extraction and interpretability. Given the significant resources already invested in projects like Gemma Scope (Lieberum et al., 2024), allocating additional compute to enhance representation quality before decoding may be justified. Importantly, more sophisticated encoders can still maintain the linear decoder necessary for downstream tasks such as steering. Empirical studies comparing feature quality across encoder complexities will be important, as will addressing concerns about non-zero-centered representations raised by Hobbhahn (2023).
@@
Our study has several limitations that warrant consideration. We primarily explored scenarios with constant sparsity and uncorrelated channels in the sparse representation, which may not fully capture the complexity of real-world data. Additionally, our analysis focused on vanilla implementations of the models, which are susceptible to issues like shrinkage due to the L1 penalty. To comprehensively understand the current amortisation gap, future work should incorporate recent SAE variants such as top-k SAEs (Makhzani & Frey, 2013; Gao et al., 2024) and JumpReLU SAEs (Rajamanoharan et al., 2024b), as discussed in Appendix F.1. This extension would allow us to analyse how large the amortisation gap is with the new SAE architectures. Similarly, our implementation of SAE+ITO did not use advanced techniques like matched pursuit or gradient pursuit, potentially underestimating its performance. The traditional dictionary learning approaches explored in Appendix F suggest room for improvement in this area. Lastly, our synthetic data generation process did not account for varying feature importance, a key aspect of Elhage et al. (2022)’s framework. Addressing these limitations in future research would provide a more comprehensive understanding of sparse encoding strategies and their applicability to complex neural representations.
@@ Appendix A Amortisation gap proof
This setting is solvable according to compressed sensing theory Donoho (2006), meaning that it is possible to uniquely recover the true $S$ up to sign flips – we cannot resolve the ambiguity between the sign of any code element and the corresponding row in the decoding matrix. If a SAE fails to achieve the same recovery, then there must be a non-zero amortisation gap, meaning that the SAE cannot solve the sparse inference problem of recovering all sparse sources from their $M$-dimensional projection. The problem is the low computational complexity of the L-NL encoder as we see by looking at its functional mapping. Essentially, the SAE is not able, not even after the nonlinear activation function, to recover the high dimensionality ($N$) of the data after a projection into a lower ($M$) dimensional space Figure 1.
@@
$\begin{split}(|s_{11}|,|s_{22}|,\gamma_{1},...,\gamma_{N-2})&=(s_{1}+s_{2})W\\ &=s_{1}W+s_{2}W\\ &=(|s_{11}|,s^{\prime}_{12},s^{\prime}_{13},...,s^{\prime}_{1N})+(s^{\prime}_{21},|s_{22}|,s^{\prime}_{23},...,s^{\prime}_{2N})\\ &=(|s_{11}|+s^{\prime}_{21},s^{\prime}_{12}+|s_{22}|,s^{\prime}_{13}+s^{\prime}_{23},...,s^{\prime}_{1N}+s^{\prime}_{2N})\end{split}$ (16)
@@
Appendix B Conceptual Modelling Framework
The concept of distributed representations in neural networks originated with the Parallel Distributed Processing (PDP) movement (Rumelhart et al., 1986). This work explored how information could be encoded across multiple units in a network, rather than in localised, symbolic representations (Thorpe, 1989). A distributed representation of information means that no single processing unit in a network performs a syntactically or semantically determinable subtask alone. Instead, an assembly of processing units generates a “distributed pattern of activation” to represent information (Smolensky et al., 2006; Van Gelder, 1990; Gelder, 1992).
We show in Figure 8 the general modelling framework we are studying for uncovering these distributed representations. Inputs pass through a neural network and generate some internal neural representation, that is often distributed and in superposition. We use an encoding process to determine active latents, or features, inherent in that process, and a learned decoder to specify what those feature directions should be. This lies at the heart of all methods studied: inference-based methods (i.e., sparse coding), amortised methods (i.e., sparse autoencoders and autoencoders with a more powerful encoder, such as an MLP) and hybrid approaches (i.e., sparse autoencoder with inference-time optimisation).
Figure 8: Illustration of Modelling Framework. Representations in neural networks (left), commonly represent input features in superposition (center left) (Elhage et al., 2022). Autoencoders can be used to extract sparse (interpretable) codes from neural representations (center right). While the decoder is fixed to be linear (an important assumption), the encoder can be more flexible. Different options for the encoder include inference, amortised inference and hybrid version of both (ITO, inference-time optimisation) (right). Moreover, the encoder might be distinct between training and testing time.
Appendix C Decoder weight analysis

### v2

@@ 4 Synthetic Sparse Inference Experiments
We present the results of our experiments comparing different sparse encoding strategies across various scenarios. All experiments were conducted using synthetic data with $N=16$ sparse sources, $M=8$ measurements, and $K=3$ active components per timestep, unless otherwise specified (more settings in Sec. 4.4, with larger values in App. B and App. C).
@@ 4.1 Known Sparse Codes
We also validated our findings at larger scales that better match real-world applications ($N=1000,M=200,K=20$, and $500,000$ data points), finding that the amortisation gap becomes even more pronounced (see Appendix C).
@@ 4.3 Unknown Sparse Codes and Dictionary
Finally, we evaluate all four methods when both latent representations and dictionary are unknown. We use a dataset of $2048$ samples, evenly split between training and testing sets, and conduct 5 independent runs of $100,000$ steps.
@@ 4.4 Performance Across Varying Data Regimes
We also investigated the trade-off between sparsity and performance for each method in Figure 6. Sparse coding achieves slightly lower reconstruction error for each L0 level, barring some very small active latents. Sparse coding shows a Pareto improvement at each L0 level in terms of MCC, even with very small active latents. The improvement is more evident when plotting against L1 rather than L0, as L1 accounts for the magnitude of non-zero values. The presence of very small non-zero latents in sparse coding motivates the exploration of top-$k$ sparse coding, detailed in Appendix H.2.
@@ 5 Interpretability of Sparse Coding Schemes
A common concern about more powerful encoding approaches is that they might learn unnatural features that are not interpretable. To investigate the interpretability of more complex encoding techniques, we trained three distinct methods on 406 million tokens from OpenWebText: a sparse autoencoder with a single linear encoder layer and ReLU activation, a multilayer perceptron encoder with one hidden layer of width 8448, and a locally competitive algorithm following the approach of Olshausen & Field (1997) and Blumensath & Davies (2008). Each method learned an overcomplete dictionary of size $16,896$ for the residual-stream pre-activations at Layer 9 of GPT-2 Small (Radford et al., 2019), which have dimension $768$.
All methods were trained using Adam with a learning rate of $3\cdot 10^{-4}$ and an $L_{1}$ penalty of $1\cdot 10^{-4}$. Following Bricken et al. (2023) and Cunningham et al. (2023), we resampled dead neurons every $15,000$ steps by setting columns with no activity to new random directions. The final results across methods were: the SAE achieved a normalised MSE of $0.061$ with a mean $L_{0}$ of 35.66 and 11% dead neurons, while the MLP reached a normalised MSE of 0.055 with a mean $L_{0}$ of $31.13$ and 22% dead neurons. The LCA approach, with $100$ gradient-based sparse-inference steps per batch, achieved a normalised MSE of 0.070. While technically none of the LCA codes were exactly zero, most were extremely small, and thresholding values below $10^{-5}$ yielded an effective $L_{0}$ of approximately $18.56$. Notably, the LCA dictionary maintained no strictly dead columns.
To assess the interpretability of the learned features, we randomly selected 500 features from each method and employed an automated interpretability classification approach using GPT-4o (full details in Appendix J). For each feature, we identified its top 10 most highly activating tokens in our 13.1 million-token test set and computed logit effects through the path expansion $W_{U}\cdot f$, where $W_{U}$ represents the model’s unembedding matrix and $f$ denotes the feature vector. We provided both the activating examples and the top and bottom 10 tokens by logit effect to GPT-4o, which generated a concise explanation of the feature’s function. To validate these interpretations, we presented them to a second instance of GPT-4o along with at least five new activating examples and five non-activating examples, labelling the activating tokens. The model predicted which examples should activate the feature based on the first instance’s explanation, allowing us to compute an F1-score against the ground truth. This automated interpretability approach is considered standard in the literature and relies on a base prompt from Juang et al. (2024).
Figure 7 displays the distributions of F1-scores across the evaluated features. The results indicate that SAE and LCA features demonstrate comparable interpretability, with median F1-scores around $0.6$. Most notably, the MLP features achieve substantially higher interpretability scores, with a median F1-score of $0.83$ and a tighter distribution. A Kruskal-Wallis test revealed significant differences between the methods ($H=1856.33$, $p<0.001$), and subsequent Dunn’s tests with Bonferroni correction confirmed that both SAE and LCA were significantly less interpretable than MLP features ($p<0.001$). See Appendix J for examples of feature interpretations.
@@
Figure 7: Distribution of F1 scores for feature interpretability across three methods (SAE, MLP, and LCA) trained on residual stream activations of Layer 9 in GPT-2. Each distribution represents 500 randomly selected features evaluated using GPT-4o for explanation generation and validation.
@@ 6 Discussion
Our study provides theoretical and empirical evidence for an inherent amortisation gap in sparse autoencoders (SAEs) when applied to neural network interpretability. We prove that SAEs with linear-nonlinear encoders cannot achieve optimal sparse recovery in settings where such recovery is theoretically possible. This limitation is supported by experimental results showing that sparse coding, and sometimes MLPs, outperform SAEs across synthetic data scenarios. Our investigation of GPT-2 activations demonstrates that MLP-based features achieve higher interpretability scores than both SAE and LCA features. These findings refute the assumption that simpler encoders are necessary for maintaining interpretability. The results carry implications for neural network interpretability, suggesting that more sophisticated encoding techniques can improve feature extraction without compromising feature validity, though at increased cost.
The use of linear-nonlinear encoders in SAEs for language model interpretability stems from concerns that more powerful methods might extract features not used by the transformer (Bricken et al., 2023). This approach appears overly restrictive given the complexity of transformer layer representations, which emerge from multiple rounds of attention and feed-forward computations. The superior performance of MLPs suggests that matching the computational complexity of the underlying representations improves feature extraction. Better encoders aligns with recent work on inference-time optimisation (Nanda et al., 2024), and will be validated as we improve encoding evaluation (Makelov et al., 2024). Regardless, SAEs are sensitive to hyperparameters and fragile (Cunningham et al., 2023), so exploring more powerful encoders is warranted.
The computational cost of complex encoders must be evaluated against gains in feature extraction and interpretability. Projects like Gemma Scope (Lieberum et al., 2024) demonstrate substantial resource investment in feature extraction, suggesting that additional compute for improved representation quality may be justified. Complex encoders can maintain the linear decoder needed for downstream tasks such as steering while providing better features. Future work should systematically compare feature quality across encoder architectures and address non-zero-centered representations (Hobbhahn, 2023).
@@
Our study has several limitations. Our LCA implementation was not optimised for the scale of experiments, requiring investigation of sparse coding methods, sparsity levels, optimisation iterations, and thresholding of near-zero activations. The gap between MLP and SAE/LCA interpretability scores warrants examination – while LCA’s lower performance likely stems from suboptimal training, the MLP’s superior interpretability relative to SAEs requires investigation. Our analysis also focused on scenarios with constant sparsity and uncorrelated channels, which may not capture real-world data complexity. Our synthetic data generation process did not account for varying feature importance described in Elhage et al. (2022)’s framework, although we did begin to explore this in Appendix D.
Future work should incorporate recent SAE variants like top-k SAEs (Makhzani & Frey, 2013; Gao et al., 2024) and JumpReLU SAEs (Rajamanoharan et al., 2024b) to measure the amortisation gap with modern architectures. Our SAE+ITO implementation did not use advanced techniques like matched pursuit, potentially underestimating its performance. The traditional dictionary learning approaches in Appendix H indicate room for improvement. Finally, we should explore sampling feature activations from different parts of the activation spectrum when doing automated interpretability, because features may exhibit different levels of specificity at different activation strengths, and examining only top activations could miss important collaborative behaviors between features and edge cases that help validate feature interpretations (Bricken et al., 2023). Addressing these limitations would advance understanding of sparse encoding strategies for complex neural representations.
@@ Appendix A Amortisation gap proof
This setting is solvable according to compressed sensing theory (Donoho, 2006), meaning that it is possible to uniquely recover the true $S$ up to sign flips – we cannot resolve the ambiguity between the sign of any code element and the corresponding row in the decoding matrix. If a SAE fails to achieve the same recovery, then there must be a non-zero amortisation gap, meaning that the SAE cannot solve the sparse inference problem of recovering all sparse sources from their $M$-dimensional projection. The problem is the low computational complexity of the L-NL encoder as we see by looking at its functional mapping. Essentially, the SAE is not able, not even after the nonlinear activation function, to recover the high dimensionality ($N$) of the data after a projection into a lower ($M$) dimensional space Figure 1.
@@
$\begin{split}(|s_{11}|,|s_{22}|,\gamma_{1},...,\gamma_{N-2})&=(s_{1}+s_{2})W\\ &=s_{1}W+s_{2}W\\ &=(|s_{11}|,s^{\prime}_{12},s^{\prime}_{13},...,s^{\prime}_{1N})\\ &+(s^{\prime}_{21},|s_{22}|,s^{\prime}_{23},\ldots,s^{\prime}_{2N})\\ &=(|s_{11}|+s^{\prime}_{21},s^{\prime}_{12}\\ &+|s_{22}|,s^{\prime}_{13}+s^{\prime}_{23},...,s^{\prime}_{1N}+s^{\prime}_{2N})\end{split}$ (14)
@@
Appendix B Relating Our Amortisation Gap to Prior Results in Sparse Autoencoders
In this appendix, we clarify how our amortisation-gap theorem aligns with prior work on shallow autoencoders in the sparse coding literature, including references such as (Rangamani et al., 2018) and (Nguyen et al., 2019). While these earlier results may appear to contradict our statement that a single feedforward linear-nonlinear encoder cannot globally recover all sparse codes from fewer measurements ($M<N$), we show that these works rely on local or probabilistic assumptions. By contrast, our theorem provides a global, worst-case statement.
Our work presents a global impossibility claim: a single-layer linear+$\sigma$ map cannot perfectly invert every $K$-sparse code if $M<N$. This argument is rank-based and does not rely on training initialisation or a specific data distribution. In contrast, many prior theorems establish local (or near-dictionary) results: they assume the encoder’s weights start sufficiently close to the true dictionary, then show that a ReLU (or threshold) gating can maintain or refine correct sparse recovery for typical data.
The distinction between uniform and distribution-specific recovery is also important. Our proof deals with uniform, adversarially chosen $K$-sparse codes. If the model must handle all codes in $\mathbb{R}^{N}$ with $\|s\|_{0}\leq K$, a single feed-forward pass will inevitably fail for some codes. By contrast, much of the prior literature – including (Rangamani et al., 2018; Nguyen et al., 2019) –requires that codes are drawn from a random distribution (e.g., sub-Gaussian or mixture-of-Gaussians). This assumption enables high-probability success on most sampled codes, but does not guarantee recovery of all codes.
Another key distinction lies in single-pass versus multi-pass inference. Our amortisation-gap statement explicitly concerns a single-layer feedforward autoencoder. Iterative or unrolled algorithms (e.g., LISTA (Gregor & LeCun, 2010), or multi-layer ReLU stacks) circumvent the rank restriction by repeatedly refining the estimate. Thus, a multi-iteration or multi-layer approach can approach near-optimal sparse recovery; but this does not contradict our statement about a one-pass linear-nonlinear encoder’s inability to decode every sparse signal.
Finally, our result demands exact (or perfect) inversion of all feasible codes, while prior analyses often accept approximate or high-probability correctness. They conclude that, given some distribution on codes and an adequately trained near-dictionary encoder, one recovers the support with probability $>1-\delta$. This does not conflict with a global impossibility statement.
Appendix C Large-Scale Experiments
To validate that our findings generalise to larger scales more representative of real-world applications, we conducted additional experiments with substantially increased dimensionality. We scaled up our synthetic experiments for the known $Z$ case to $N=1000$ sparse sources, $M=200$ measurements, and $K=20$ active components, training on $500,000$ samples for $20,000$ steps. This represents a significant increase from our base experiments (which used $N=16$, $M=8$, $K=3$), bringing us closer to the scale of actual SAE applications.
For these experiments, we modified our training procedure to use minibatch processing (batch size $1024$) to handle the increased data scale efficiently. We evaluated MLPs with hidden layer widths of $H=\{256,512,1024\}$ against a standard SAE. The results, shown in Figures 8(a) and 8(b), demonstrate that our key findings about the amortisation gap not only hold but become more pronounced at larger scales.
(a) MCC vs. training steps
(b) MCC vs. total FLOPs
Figure 8: (Larger $N$, $M$ and $K$) Performance comparison of SAE and MLPs in predicting known latent representations. The black dashed line in (b) indicates the average FLOPs at which MLPs surpass SAE performance.
Specifically, the performance gap is slightly more substantial than in our smaller-scale experiments, suggesting that the limitations of linear-nonlinear encoders become more significant as the problem dimensionality increases. This aligns with our theoretical predictions, as the higher-dimensional setting creates more opportunities for interference between features that the simple SAE encoder struggles to disentangle.
The FLOP analysis (Figure 8(b)) reveals that all MLPs surpass the SAE’s performance at approximately $3\times 10^{14}$ FLOPs, regardless of hidden layer width. This consistent computational threshold, despite varying model capacities, suggests a fundamental limitation in the SAE’s architecture rather than a simple capacity constraint.
Appendix D A Different Distribution of Codes
In this appendix, we explore an alternative data generation process that better reflects the distributional properties observed in real-world latent representations. While our main experiments use uniformly sampled sparse codes, recent work has shown that latent features in large models often follow heavy-tailed distributions (e.g., power laws) with varying activation frequencies (Engels et al., 2024; Park et al., 2024). To investigate the robustness of our findings, we modify our synthetic data generator to incorporate a Zipf distribution (parameterised by $\alpha$) over feature activations. This creates a hierarchical structure where certain features are consistently more likely to be active and have larger magnitudes, while others are more rarely activated. The modified generator maintains the core sparsity constraint of $K$ active dimensions, but weighs both the selection probability and magnitude of each dimension according to its position in the Zipf distribution.
(a) MCC vs. training steps
(b) MCC vs. total FLOPs
Figure 9: (Zipfian) Performance comparison of SAE and MLPs in predicting known latent representations. The black dashed line in (b) indicates the average FLOPs at which MLPs surpass SAE performance.
We reproduced all experiments from Section 4.4 using this modified data generation process, with $\alpha=1.0$. The results reveal several interesting differences while broadly supporting our main conclusions. In the known sparse codes scenario (Figure 9), all methods achieve higher absolute performance, with MLPs reaching MCC values of approximately 0.8 compared to 0.6 in the uniform case. The advantage of wider hidden layers becomes more pronounced under the Zipfian distribution, though the computational threshold at which MLPs surpass SAE performance remains consistent with our original findings.
(a) MCC vs. training steps
(b) MCC vs. total FLOPs
Figure 10: (Zipfian) Performance comparison of SAE, SAE with inference-time optimisation (SAE+ITO), and MLPs in predicting latent representations with a known dictionary. Dashed lines in (b) indicate extrapolated performance beyond the measured range.
When the dictionary is known but sparse codes are unknown (Figure 10), we observe similar relative performance patterns but with higher peak MCC values (around 0.85 compared to 0.75 in the uniform case). The SAE with inference-time optimisation (SAE+ITO) exhibits more volatile training dynamics under the Zipfian distribution, showing a characteristic performance drop around $10^{4}$ training steps before recovery. This suggests that optimisation becomes more challenging when dealing with hierarchically structured features, though the method ultimately achieves strong performance.
(a) Latent prediction: MCC vs. training steps
(b) Latent prediction: MCC vs. total FLOPs
(c) Dictionary learning: MCC vs. training steps
(d) Dictionary learning: MCC vs. total FLOPs
Figure 11: (Zipfian) Dictionary learning performance comparison when both $s^{*}$ and $D^{*}$ are unknown.
The most substantial differences emerge in the fully unsupervised setting, where both dictionary and sparse codes are unknown (Figure 11). Here, the Zipfian distribution leads to lower overall performance (MCC of 0.5-0.6 versus 0.7-0.8 in the uniform case) and creates clearer separation between different methods. While sparse coding still outperforms other approaches, its advantage is less pronounced than in the uniform setting. Dictionary learning under the Zipfian distribution shows increased volatility across all methods, particularly for sparse coding, though the relative ordering of performance remains consistent with our original results.
These findings suggest that while our conclusions about the relative merits of different approaches hold under more realistic distributional assumptions, the absolute difficulty of the sparse inference problem increases when dealing with hierarchically structured features.
Appendix E Decoder weight analysis